Grade 9General Physics

A person sitting on the top of a tall building is dropping ball at regular intervals of one second .find the positions of the 3rd ,4th and 5 th ball when the 6th ball is being dropped

Aneesh
9 Years agoGrade 9

1 Answer

Vikas TU
9 Years ago
Dear Student,

when the 6th ball is dropped, the 5th ball must have travelled for one second, the 4th for 2 seconds and the third one for 3 seconds.

initial velocity for each ball is 0 since they are dropped.

S=ut+at2/2

Distance travelled by the 5th ball=g/2=5m
Distance travelled by the 4th ball=4g/2=20m
Distance travelled by the 3rd ball= 9g/2 = 45m.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)